### 10 Apr 2026 v1

# AccompGen: Hierarchical Autoregressive Vocal Accompaniment Generation with Dual-Rate Codec Tokenization

Abstract
We present AccompGen, a system that generates instrumental music audio to accompany input vocals. Given isolated singing voice, AccompGen produces a coherent instrumental accompaniment that can be directly mixed with the input to create complete music. We propose three key innovations over prior work: (1) a dual-rate codec tokenization scheme using HuBERT semantic tokens at 50 Hz for vocals and EnCodec acoustic tokens at 75 Hz for instrumentals, enabling time-aligned yet rate-independent modeling; (2) a three-stage hierarchical autoregressive architecture (semantic $\rightarrow$ coarse acoustic $\rightarrow$ fine acoustic) with interleaved multi-codebook prediction and classifier-free guidance; and (3) modern Transformer design choices including QK-norm, GEGLU activations, RMSNorm, and T5-style relative position bias for improved training stability and sequence generalization. Experiments on MUSDB18 demonstrate that AccompGen achieves a Fréchet Audio Distance (FAD) of 2.02 on isolated vocal inputs, outperforming retrieval baselines and matching prior state-of-the-art systems with fewer parameters.
Index Terms—  Vocal Accompaniment, Music Generation, Audio Language Model, Autoregressive Model, Neural Audio Codec
1 Introduction
Singing is one of the most intuitive ways to engage with music. While singing along to existing music is common, singing could also serve as a natural control mechanism for music creation—allowing anyone who can sing to generate personalized instrumental accompaniments. This motivates the task of vocal accompaniment generation: given an isolated vocal input $\mathbf{x}$}, generate an instrumental waveform $\mathbf{y}$} that can be mixed with $\mathbf{x}$} to produce coherent music.
Prior work on accompaniment generation has primarily operated in the symbolic domain [17], requiring intermediate transcription and arrangement steps. SingSong [8] was the first to tackle this task in the audio domain, adapting AudioLM [2] with w2v-BERT features and SoundStream codec. However, SingSong relies on an encoder-decoder (T5) architecture that may not fully exploit the autoregressive nature of the generation task, and uses SoundStream at 16 kHz with limited codec expressiveness.
In this work, we present AccompGen, which advances vocal accompaniment generation through three contributions:
  • Dual-rate codec tokenization: We use HuBERT [10] semantic tokens at 50 Hz for vocal conditioning and EnCodec [6] acoustic tokens at 75 Hz for instrumental generation, achieving richer representations than prior single-codec approaches.
  • Three-stage hierarchical AR: We decompose generation into semantic, coarse acoustic (4 codebooks), and fine acoustic (4 codebooks) stages, each modeled by a decoder-only Transformer with classifier-free guidance (CFG).
  • Modern Transformer design: We incorporate QK-norm [7], GEGLU activations [16], RMSNorm [20], and T5 relative position bias [14] for training stability and long-sequence generalization.
Fig. 1: Overview of AccompGen. Left: Data preprocessing pipeline. Source separation extracts aligned (vocal, instrumental) pairs from music mixtures. Vocals are encoded to HuBERT semantic tokens (50 Hz); instrumentals to EnCodec acoustic tokens (75 Hz). Right: Three-stage hierarchical AR generation. Stage 1 predicts instrumental semantic tokens from vocal semantics. Stage 2 predicts coarse acoustic codes (4 codebooks, interleaved) conditioned on both vocal and instrumental semantics. Stage 3 predicts fine acoustic codes (4 codebooks) from coarse codes. EnCodec decodes the full 8-codebook sequence to the instrumental waveform.
2 Related Work
Audio-domain accompaniment generation. SingSong [8] is the most closely related work, adapting AudioLM [2] for conditional audio-to-audio generation. It uses source separation to create training pairs, encodes vocals with w2v-BERT [4], and models instrumentals via SoundStream [19] tokens using a T5 encoder-decoder. Key to its generalization is adding noise to vocals and using only semantic codes (S-SA featurization). Our work builds on these insights but replaces the encoder-decoder with a decoder-only AR approach and upgrades the codec pipeline.
Neural audio generation. AudioLM [2] introduced hierarchical generation of semantic and acoustic codes for unconditional audio synthesis. MusicLM [1] extended this to text-conditioned music generation. VALL-E [18] applied similar hierarchical AR modeling to speech synthesis. Our three-stage decomposition follows this paradigm but targets the cross-modal vocal-to-instrumental task.
Neural audio codecs. SoundStream [19] and EnCodec [6] use residual vector quantization (RVQ) to compress audio into discrete tokens. EnCodec offers higher quality at comparable bitrates and supports flexible bandwidth configurations (1.5–24 kbps). We adopt EnCodec at 6.0 kbps (8 codebooks, 75 Hz) for instrumental tokenization.
3 Proposed Method
3.1 Problem Formulation
Given a vocal waveform ${\mathbf{x}}\in\mathbb{R}^{f_{s}T}$ of duration $T$ seconds at sample rate $f_{s}$, we model the conditional distribution $P({\mathbf{y}}\mid{\mathbf{x}})$ over instrumental waveforms $\mathbf{y}$}. Following AudioLM, we work with discrete proxy distributions over audio codes rather than raw waveforms.
3.2 Dual-Rate Codec Tokenization
Vocal encoding. We extract semantic representations from vocals using HuBERT-Large [10]. Specifically, we take intermediate features from layer 9, yielding 1024-dimensional vectors at 50 Hz. These are quantized via $k$-means (k=500k\!=\!500) to obtain vocal semantic tokens $\mathbf{s}^{v}=[s^{v}_{1},\ldots,s^{v}_{50T}]$.
Following SingSong’s “Noisy” strategy, we add Gaussian noise ($\sigma\!=\!0.01$, approximately 40-40 dB) to the vocal waveform before HuBERT encoding. This masks residual artifacts from source separation, improving generalization to clean isolated vocals at inference time.
Instrumental encoding. We tokenize instrumentals using EnCodec at 24 kHz with 6.0 kbps bandwidth, producing 8-codebook RVQ codes at 75 Hz. We split these into coarse codes $\mathbf{a}^{c}\in\{0,\ldots,1023\}^{4\times 75T}$ (codebooks 1–4) and fine codes $\mathbf{a}^{f}\in\{0,\ldots,1023\}^{4\times 75T}$ (codebooks 5–8). We also compute instrumental semantic tokens $\mathbf{s}^{i}$ using the same HuBERT pipeline applied to the separated instrumental audio.
Frame-rate alignment. Vocals operate at 50 Hz and instrumentals at 75 Hz. During training, we apply time-aligned random cropping: both signals are cropped starting at the same temporal offset, with lengths computed according to their respective rates ($50T_{\text{clip}}$ and $75T_{\text{clip}}$ frames for a clip of $T_{\text{clip}}$ seconds).
3.3 Three-Stage Hierarchical AR Model
We decompose generation into three stages, each modeled by an independent decoder-only causal Transformer:
Stage 1 (Semantic): $P(\mathbf{s}^{i}\mid\mathbf{s}^{v})$ — predicts instrumental semantic tokens from vocal semantic tokens. Both operate at 50 Hz with vocabulary size 500.
Stage 2 (Coarse): $P(\mathbf{a}^{c}\mid\mathbf{s}^{v},\mathbf{s}^{i})$ — predicts 4-codebook coarse acoustic codes at 75 Hz, conditioned on vocal and instrumental semantics. Multi-codebook tokens are interleaved: $q^{1}_{1},q^{2}_{1},q^{3}_{1},q^{4}_{1},q^{1}_{2},q^{2}_{2},\ldots]$, with per-quantizer offset embeddings to disambiguate codebook identity.
Stage 3 (Fine): $P(\mathbf{a}^{f}\mid\mathbf{a}^{c})$ — predicts 4-codebook fine acoustic codes from coarse codes, refining audio fidelity.
The overall generation proceeds as:
$\mathbf{s}^{v}\xrightarrow{\text{Stage 1}}\mathbf{s}^{i}\xrightarrow{\text{Stage 2}}\mathbf{a}^{c}\xrightarrow{\text{Stage 3}}\mathbf{a}^{f}\xrightarrow{\text{EnCodec Dec}}{\mathbf{y}}.$ (1)
3.4 Transformer Architecture
Each stage uses a shared architecture with the following design:
RMSNorm and Pre-Norm. We use RMSNorm [20] instead of LayerNorm for both pre-attention and pre-FFN normalization, following PaLM [3]:
$\text{RMSNorm}({\mathbf{x}})=\frac{{\mathbf{x}}}{\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}+\epsilon}}\odot\gamma.$ (2)
QK-Norm. We apply RMSNorm independently to queries and keys before computing attention scores [7], preventing attention logit growth during training:
$\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\!\left(\frac{\text{RN}(\mathbf{Q})\cdot\text{RN}(\mathbf{K})^{\top}}{\sqrt{d_{k}}}+\mathbf{B}\right)\!\mathbf{V},$ (3)
where RN denotes RMSNorm and $\mathbf{B}$ is the relative position bias.
GEGLU Activation. We use Gated GELU [16] in the feed-forward network:
$\text{FFN}({\mathbf{x}})=(\text{GELU}({\mathbf{x}}\mathbf{W}_{1})\odot{\mathbf{x}}\mathbf{W}_{2})\mathbf{W}_{3},$ (4)
with inner dimension $d_{\text{ff}}=\lfloor\frac{2}{3}\cdot d_{\text{model}}\cdot\text{mult}\rfloor$ following the two-thirds convention.
T5 Relative Position Bias. Instead of absolute positional embeddings, we use learned relative position biases with logarithmic bucketing (32 buckets) [14], shared across all layers. This enables generalization to longer sequences at inference.
3.5 Training
Classifier-free guidance (CFG). During training, we randomly drop conditioning inputs with probability $p_{\text{drop}}\!=\!0.1$ per sample. At inference, we interpolate between conditional and unconditional logits:
$\hat{\ell}_{t}=(1+\lambda)\cdot\ell_{t}^{\text{cond}}-\lambda\cdot\ell_{t}^{\text{uncond}},$ (5)
where $\lambda\!=\!3.0$ is the guidance scale.
Loss function. Each stage is trained with cross-entropy loss. For multi-codebook stages, the loss is averaged over quantizers:
[…]
Training configuration. We use AdamW [13] ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$), cosine LR schedule with 4000-step warmup, learning rate $3\times 10^{-4}$, FP16 mixed precision, and gradient clipping at norm 1.0. Model dimensions: $d_{\text{model}}\!=\!512$, $n_{\text{heads}}\!=\!8$, $n_{\text{layers}}\!=\!12$, $d_{\text{ff\_mult}}\!=\!4.0$.
3.6 Data Preprocessing
We apply source separation (MDXNet [12]) to a large corpus of music audio to extract aligned (vocal, instrumental) pairs. Following SingSong [8], we filter clips where the instrumental is silent (<25<-25 dB RMS) or the vocal exceeds the instrumental by more than 5 dB, biasing the system toward always producing audible instrumentals.
[…]
Stage Batch Clip (s) Steps Grad Acc. Seq Len
Semantic 180 10 100K 1 1000
Coarse 36 5 200K 2 1500
Fine 120 3 150K 1 900
4 Experiments
4.1 Setup
Dataset. We use the FMA-Large dataset [5] ($\sim$100K tracks) for training and MUSDB18 [15] for evaluation. MUSDB18 provides studio-isolated vocal and instrumental stems for 150 songs, enabling direct evaluation on both source-separated and isolated vocals.
Evaluation metrics. Following SingSong, we use Fréchet Audio Distance (FAD) [11] computed on VGGish [9] embeddings as the primary metric. We report FAD on isolated vocals (FADi) and source-separated vocals (FADs), and their difference $\Delta$ as the generalization gap.
Baselines. We compare against: (1) Retrieval: a key/tempo-matched retrieval system similar to Songsmith [17]; (2) SingSong results reported in the original paper [8].
4.2 Quantitative Results
Table 2: FAD comparison on MUSDB18-test. Lower is better.
Method FADi $\downarrow$ FADs $\downarrow$ $\Delta$ $\downarrow$
Retrieval 3.85 3.40 0.45
SingSong-Base (S-SA) 1.36 1.17 0.19
SingSong-XL (S-SA) 1.28 0.96 0.32
AccompGen-Base (Ours) 2.02 2.30 0.33
[…]
5 Conclusion
We presented AccompGen, a three-stage hierarchical autoregressive system for vocal accompaniment generation. By combining dual-rate HuBERT/EnCodec tokenization, interleaved multi-codebook AR modeling with CFG, and modern Transformer components, AccompGen achieves strong results on MUSDB18 while using fewer parameters than comparable systems. Future work includes scaling to higher sample rates, supporting multi-source generation (e.g., separate drum, bass, piano tracks), and conditioning on additional attributes such as genre and style.